AnxiTEA:

In a bag full of small balls, 1/4 of these balls are green, 1/8 are blue, 1/12 are yellow and the remaining 26 white. How many balls are blue?

MxxnLight:

Any idea?

AnxiTEA:

i dont have any clue, my brain isn't working

Angle:

how much is: 1/4 + 1/8 + 1/12 = ?

AnxiTEA:

idk ;-;

XioGonz:

X= balls x/4 green = 6 x/24 x/8 blue = 3 x/24 x/12 yellow = 2 x/24 sum = 11 x/24 13 x/24 are white 13 x/24 = 26 x = 24 * 26/13 x/8 = blue = 3 *26/13 = 3*2 = 6

AnxiTEA:

6?

AnxiTEA:

oou Thank you!

AnxiTEA:

:D